OpenStudy (anonymous):

find the fraction form of 3.02. 2 is recuring. please help ASAP

OpenStudy (anonymous):

you cant have 2 decimal points in a number...

OpenStudy (anonymous):

its not 2 decimal points. 3.02 2 is recuring

OpenStudy (anonymous):

this is from sequence and series module

OpenStudy (anonymous):

oh sorry nvm then :\ wish i could help srry bro

OpenStudy (anonymous):

hmm i am guessing you may have to correct me though but is it 302 -3.02 298.98

OpenStudy (anonymous):

this is what i got

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if thats the recurring num then wrong answer

OpenStudy (anonymous):

yea and its from sequence and series module do you know about that

OpenStudy (anonymous):

show me your work please

OpenStudy (anonymous):

lol i think this is memorized

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

did you delete it why?

OpenStudy (anonymous):

please help guys

OpenStudy (anonymous):

my connection is playing up

OpenStudy (anonymous):

tryin to convert 3.02222222222... to fraction?

OpenStudy (anonymous):

umm i still getting confused it is still somwhere between my first answer

OpenStudy (anonymous):

x=3.02 10x=30.2 100x=302.2 -(10x=30.2) 90x=272 \[x =\frac{272}{90} = \frac{136}{45} \]